The common stock of Flavorful Teas has a beta of 48. The risk-free rate of return...

  1. The common stock of Flavorful Teas has a beta of 48. The risk-free rate of return is 2.5 percent.
  1. The return on the market is 12 percent and. What is the expected return on this stock?
  2. If the Market risk premium is 8.5%, what is the expected return on this stock?

  1. ABC Corp just paid a dividend of $1.5 per share. The dividend is expected to grow at 4% a year indefinitely. If the required rate of return is 8.5%, wat is the current value of ABC Corp?

  1. The Bell Weather Co. is a new firm in a rapidly growing industry. The company is planning on increasing its annual dividend by 12% a year for the next four years and then decreasing the growth rate to 5 percent per year. The company just paid its annual dividend in the amount of $1.5 per share. What is the current value of one share of this stock if the required rate of return is 10 percent?
Answer #1

Ques 1). Given for Flavorful Teas

Beta = 48 (assuming this beta given is correct. if not you can substitute its new value in the equation)

risk free rate Rf = 2.5%

a). Return on market Rm = 12%

So, expected return on stock using CAPM is

Ke = Rf + Beta*(Rm - Rf) = 2.5 + 48*(12-2.5) = 458.5%

b). When market risk premium MRP = 8.5%

So, expected return on stock using CAPM is

Ke = Rf + Beta*(Rm - Rf) = 2.5 + 48*8.5 = 410.5%

Ques 2). Given for ABC corp

Current dividend D0 = $1.5

Growth rate g = 4%

required rate of return Ke = 8.5%

So current stock price using constant dividend growth model is

P0 = D0*(1+g)/(Ke-g) = 1.5*1.04/(0.085-0.04) = $34.67

Request you to post other question separately, as HomeworkLib policy does not allow this.